OpenStudy (anonymous):

Which of the following are solutions to the equation below? Check all that apply. 3x2 + 27x + 54 = 0 A. -6 B. -3 C. 6 D. 3 E. 9 F. -27

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

tomake it a little easier you can divide each term by 3

OpenStudy (welshfella):

giving x^2 + 9x + 18

OpenStudy (welshfella):

can you thinl of 2 numbers whose product is 18 and sum is 9?

OpenStudy (anonymous):

i have to find what goes into 27 and 54?

OpenStudy (welshfella):

- iv dodne that for you to give an easier equation

OpenStudy (anonymous):

what is product?

OpenStudy (anonymous):

=?

OpenStudy (anonymous):

3*6 3+6

OpenStudy (welshfella):

product mean result when 2 nubers are multiplied like 6 is the product of 2 an 3

OpenStudy (welshfella):

Yea you got it 3 and 6 are the 2 numbers we want

OpenStudy (welshfella):

so factors are (x + 3(x + 6) = 0 so the roots are what? remember what we did for the other equation

OpenStudy (anonymous):

sub the opposite?

OpenStudy (welshfella):

x+ 3 = 0 or x + 6 = 0 so x = what? or what?

OpenStudy (welshfella):

x + 3 = 0 x + 3 -3 = 0 - 3

OpenStudy (welshfella):

so x = ?

OpenStudy (anonymous):

-3 -6?

OpenStudy (welshfella):

-3 and -6 yes

OpenStudy (anonymous):

thank you <3

OpenStudy (anonymous):

will it ever be positive?

OpenStudy (welshfella):

yw I think you are understanding it a bit more now

OpenStudy (welshfella):

yea if for example x - 3 = 0 then x - 3 + 3 = 0 + 3 x = 3

OpenStudy (anonymous):

oh okay